OpenStudy (anonymous):

Show that \[\frac{\sqrt{x^2 + x}}{x} = \sqrt{\frac{1}{x} + 1}\] I'm not sure how to go about the algebraic manipulation to turn it into that result.

OpenStudy (anonymous):

You see the x on the left hand side? In order to bring that under the square root above, you need to square it (so that its square root will take you back to x again), then just divide through...

OpenStudy (anonymous):

So then I get: \[\frac{\sqrt{\sqrt{x^2 + x}^2}}{\sqrt{x^2}}\] Is this right?

OpenStudy (anonymous):

It's just sqrt(x^2 + x/x^2)

OpenStudy (anonymous):

Yea, but doesn't squaring the terms change the value of the expression?

OpenStudy (anonymous):

You start with x on the bottom and bring it under the square root as x^2 This is just a rearrangement, x is same as sqrt(x^2)

OpenStudy (anonymous):

Oh, so I don't have to square then square root again for the numerator, only for the denominator

OpenStudy (anonymous):

\[\frac{ \sqrt{x ^{2}+x} }{ \sqrt{x ^{2}} }\]

OpenStudy (anonymous):

Right, you are only "moving" the x "into" the square root, so that is all you need to square

OpenStudy (anonymous):

Then you can rewrite like\[\sqrt{\frac{ x ^{2}+x }{ x ^{2} }}\]Can you simplify from there?

OpenStudy (anonymous):

Yea, if I knew I could do that it's easy but I didn't know you were allowed to do that in algebra

OpenStudy (anonymous):

As, you said, the key thing is not to change the value, as long as you don't do that, no problem

OpenStudy (anonymous):

its just manipulation, making the equation look different without really changing anything. Like @estudier said, you can change the x in the denominator to \[\sqrt{x ^{2}}\]because that is just x written differently

OpenStudy (anonymous):

you can think of it in reverse, if I give you sqrt(a^2/b^2) you can bring out the b^2 as b -> sqrt(a^2)/b